# GHOST 2.0: Generative High-fidelity One Shot Transfer of Heads

### v3

@@ 1 Introduction
Currently there are only few papers that address the task of head swap. The first steps were taken by DeepFaceLab (Perov et al., 2020), which requires finetuning for each source identity. Next work, StylePoseGAN (Sarkar et al., 2021), tends to modify background and skin color of the target. A seminal work that partially tackles the aforementioned issues is HeSer (Shu et al., 2022). It splits the generation process into head reenactment stage and reference-based blending stage. This work was subsequently extended to a diffusion setting by FaceX model (Han et al., 2023). However, both methods still face issues concerning the quality of head generation and color transfer.
@@ 2 Related Work
The task of head swap is covered by a limited number of works. DeepFaceLab (Perov et al., 2020) is the first approach enabling this capability. However, it requires a large amount of source data for training, and poorly performs color transfer and fusion of generated head with the background. StylePoseGAN (Sarkar et al., 2021) performs head swap by conditioning StyleGAN (Karras et al., 2020) on pose and combined texture map, with body parts taken from target and head — from source. Still, it tends to modify the background and skin color of the target. HeSer (Shu et al., 2022) tackles these issues by designing a separate module for each task. First, it uses head reenactment model based on (Burkov et al., 2020) to align source head with target in pose and expression. In the second stage a reference is created for skin color transfer based on correlation between pixels from the same head parts. Together with background inpainting prior, it is used to condition blending UNet (Ronneberger et al., 2015) which fuses the head with background. While this method outperforms the competitors, it suffers from identity leakage of target and is unable to color head parts present in source image but absent in the driving one. While previous methods are based on GANs (Goodfellow et al., 2014), there have been attempts (Wang et al., 2022a; Han et al., 2023) to use diffusion models (Ho et al., 2020) instead. However, currently these approaches face issues of pose controllability, preservation of target skin color and overall realism of generated head.
@@ 3.2.1 Preliminary
When transferring to real data, we must take into account that our model works within cropped images. In contrast to the face swap problem, where the person’s face is always inside the generated region, in the head swap problem the hair can extend beyond the boundaries of the considered area. In this case, we need to remove excess hair from the image. For this purpose, we propose a post-processing blending algorithm using Kandinsky diffusion model (Shakhmatov et al., 2023). We obtain the final image by inpaiting the region masked by $M_{inpainting}$ for UNet output $I_{B}$, as shown in fig. 8. It should be noted that this step is optional and is applied when the person in the driving image has long hair.
@@ 4.3 Blender evaluation
To evaluate our Blender and justify external inpainting usage in it, we trained baseline HeSer Blender on the same data. As can be seen from table 4, version with external inpainting significantly outperforms baseline version on every metric. Our solution also doesn’t fail in hard cases when source and target differ significantly (fig. 11).
@@ 5 Conclusion
The limitations of our model are the following. In several cases, our method does not reproduce fine details of source appearance. This can be tackled by using stronger appearance encoders in Aligner. Concerning Blender, some face parts may not be evenly colored if the area of corresponding color reference is small. Solving these problems constitutes a direction for future work.

### v4

@@ 1 Introduction
Currently there are only few papers that address the task of head swap. The first steps were taken by DeepFaceLab (Perov et al., 2020), which requires finetuning for each source identity. Next work, StylePoseGAN (Sarkar et al., 2021), tends to modify background and skin color of the target. A seminal work that partially tackles the aforementioned issues is HeSer (Shu et al., 2022). It splits the generation process into head reenactment stage and reference-based blending stage. This work was followed by diffusion-based approaches (Baliah et al., 2025; Han et al., 2023). However, latest methods still face issues concerning quality of head generation and color transfer.
@@ 2 Related Work
The task of head swap is covered by a limited number of works. DeepFaceLab (Perov et al., 2020) is the first approach enabling this capability. However, it requires a large amount of source data for training, and poorly performs color transfer and fusion of generated head with the background. StylePoseGAN (Sarkar et al., 2021) performs head swap by conditioning StyleGAN (Karras et al., 2020) on pose and combined texture map, with body parts taken from target and head — from source. Still, it tends to modify the background and skin color of the target. HeSer (Shu et al., 2022) tackles these issues by designing a separate module for each task. First, it uses head reenactment model based on (Burkov et al., 2020) to align source head with target in pose and expression. In the second stage a reference is created for skin color transfer based on correlation between pixels from the same head parts. Together with background inpainting prior, it is used to condition blending UNet (Ronneberger et al., 2015) which fuses the head with background. While this method outperforms the competitors, it suffers from identity leakage of target and is unable to color head parts present in source image but absent in the driving one. While previous methods are based on GANs (Goodfellow et al., 2014), there have been attempts (Baliah et al., 2025; Han et al., 2023; Wang et al., 2022a) to use diffusion models instead (Ho et al., 2020). However, these approaches face issues of pose controllability, preservation of target skin color and overall realism of generated head.
@@ 3.2.1 Preliminary
When transferring to real data, we must take into account that our model works within cropped images. In contrast to the face swap problem, where the person’s face is always inside the generated region, in the head swap problem the hair can extend beyond the boundaries of the considered area. In this case, we need to remove excess hair from the image. For this purpose, we propose a post-processing blending algorithm using Kandinsky diffusion model (Shakhmatov et al., 2023). We obtain the final image by inpaiting the region masked by $M_{inpainting}$ for UNet output $I_{B}$, as shown in fig. 8. It should be noted that this step is optional and is applied when person in the target image has hair that significantly differs from hair of source.
@@ 4.3 Blender evaluation
To evaluate our Blender and justify external inpainting usage in it, we trained baseline HeSer Blender on the same data. Additionally, we compare with diffusion-based head swap approaches, such as REFace (Baliah et al., 2025) and InstantID (Wang et al., 2024). For the latter solution, we first use InstantID (Wang et al., 2024) to generate source head with target pose directed by facial keypoints. Then the resulting head is blended into target background using SDXL (Podell et al., 2023) inpainting model. It should be noted that both solutions frequently fail to preserve identity and target skin color. As can be seen from table 4, our version with external inpainting outperforms baseline and competitors on almost every metric. However, for background inpainting, SDXL shows slightly better results than LaMa, so further investigation into better inpainting models may constitute an area for future research. Furthermore, for external inpainting and baseline, we present results on training with additional data augmentations. Our solution outperforms baseline in this scenario as well and also doesn’t fail in hard cases when source and target differ significantly (fig. 11).
In addition, we present inference results of our model on real-life and outdoor photos in supplementary 17. Although our model successfully copes with most cases, sometimes Kandinsky model hallucinates during post-processing and may produce additional attributes such as collars when inpainting excess hair. Also, hair for reenacted source head is generated only inside the crop region. In case target person has shorter hair, it is likely that in the resulting image source hair would look unnaturally cropped from below.
@@ 5 Conclusion
The limitations of our model are the following. In several cases, our method does not reproduce fine details of source appearance. This can be tackled by using stronger appearance encoders in Aligner. Concerning Blender, some face parts may not be evenly colored if the area of corresponding color reference is small. Additionally, blending stage could be improved by mitigating the generation of additional clothing attributes (e.g. collars) and better hair processing. Complex lighting conditions can also produce inferior results during blending.
@@ Appendix D Further ablations on Aligner
Figure 15: Head swap results with the same identity and color augmentation applied to the source
Figure 16: Head swap results with different identities
Figure 17: Head swap results on real-life and outdoor photos